# TARDis: Time Attenuated Representation Disentanglement for Incomplete Multi-Modal Tumor Segmentation and Classification

### v1

@@ Abstract
Tumor segmentation and diagnosis in contrast-enhanced Computed Tomography (CT) rely heavily on the physiological dynamics of contrast agents. However, obtaining a complete multi-phase series is often clinically unfeasible due to radiation concerns or scanning limitations, leading to the "missing modality" problem. Existing deep learning approaches typically treat missing phases as absent independent channels, ignoring the inherent temporal continuity of hemodynamics. In this work, we propose Time Attenuated Representation Disentanglement (TARDis), a novel physics-aware framework that redefines missing modalities as missing sample points on a continuous Time-Attenuation Curve. TARDis explicitly disentangles the latent feature space into a time-invariant static component (anatomy) and a time-dependent dynamic component (perfusion). We achieve this via a dual-path architecture: a quantization-based path using a learnable embedding dictionary to extract consistent anatomical structures, and a probabilistic path using a Conditional Variational Autoencoder to model dynamic enhancement conditioned on the estimated scan time. This design allows the network to hallucinate missing hemodynamic features by sampling from the learned latent distribution. Extensive experiments on a large-scale private abdominal CT dataset (2,282 cases) and two public datasets demonstrate that TARDis significantly outperforms state-of-the-art incomplete modality frameworks. Notably, our method maintains robust diagnostic performance even in extreme data-sparsity scenarios, highlighting its potential for reducing radiation exposure while maintaining diagnostic precision.
@@ 1 Introduction
Tumor segmentation in contrast-enhanced Computed Tomography (CT) is fundamentally dependent on the transient physiological dynamics of the contrast agent [bi2019artificial]. Crucially, the diagnostic boundaries of tumors are often defined by this time-dependent contrast-enhancement signal, which is unavailable in non-contrast acquisitions or incorrectly timed scans [bae2010intravenous]. Accurately capturing these dynamics usually requires a multi-phase acquisition protocol. However, obtaining a complete set of phases is not always feasible in clinical practice due to radiation dose concerns, patient motion, or limited scanning time. This creates a significant "missing modality" problem, where crucial enhancement information is absent, leading to severe degradation in segmentation performance for standard learning-based frameworks [javed2024robustness].
@@
Existing approaches to multi-modal segmentation typically rely on channel-concatenation strategies, which assume a fixed set of inputs. When a modality is missing, these models often fail or require retraining. To address this, recent "incomplete modality" frameworks have been proposed, utilizing techniques such as arithmetic mean fusion, knowledge distillation, or shared latent space mapping [dou2020unpaired, chen2019robust]. While these methods allow for flexible inputs, they generally treat different contrast phases as independent, uncorrelated channels or simple statistical variations. By ignoring the temporal dependency between phases, existing methods struggle to infer the missing dynamic information effectively, often resulting in suboptimal feature representation [shen2019brain, sharma2019missing].
@@
In this work, we propose Time Attenuated Representation Disentanglement (TARDis), a novel framework that addresses the incomplete modality challenge by modeling the underlying hemodynamics of CT. Our central hypothesis is that the Hounsfield Unit (HU) intensity can be explicitly decoupled into a time-invariant static component and a time-dependent dynamic component. Rather than treating missing modalities as empty data channels, we treat them as missing samples on the TAC. Our method employs a shared encoder to process a flexible number of input modalities. To learn robust representations, we introduce a dual-path disentanglement mechanism. First, we extract a modal-agnostic anatomical representation using a learnable embedding dictionary, ensuring consistency across different scans. Second, we model the modal-specific dynamic representation using a Conditional Variational Autoencoder (CVAE) [sohn2015learning], which is conditioned on the anatomy and the estimated scan time. This allows the network to probabilistically reconstruct the dynamic enhancement features of the tumor even when specific phases are missing. By explicitly disentangling anatomy from time-dependent perfusion, TARDis learns a physically grounded latent space that is robust to missing data, superior to standard fusion or masking strategies.
@@ 2.1 Multi-Modal Medical Image Segmentation
Multi-modal learning has become a cornerstone in medical image analysis, as combining complementary information from different modalities significantly improves diagnostic precision [9363915]. Early approaches focused on input-level fusion, where registered modalities are concatenated as multi-channel inputs to a Convolutional Neural Network [kamnitsas2017efficient, milletari2016v]. While effective, this assumes a fixed set of inputs, making the model fragile to missing data. Feature-level fusion strategies have subsequently emerged to better exploit cross-modal correlations. For instance, Dolz et al. [dolz2018hyperdense] proposed HyperDenseNet to facilitate dense connectivity across multi-modal streams. More recently, Transformer-based architectures have been adopted to model long-range dependencies across modalities. Hatamizadeh et al. [hatamizadeh2022unetr] introduced UNETR, which utilizes a vision transformer encoder to fuse volumetric data. Despite their success, these "complete-modality" frameworks rely on the assumption that all training and testing modalities are available. In clinical practice—particularly in multi-phase CT—scans are often incomplete due to radiation dose concerns or varying acquisition protocols, rendering these rigid fusion models ineffective without retraining or zero-filling strategies that often introduce artifacts.
@@ 2.2 Learning with Incomplete Modalities
To address the "missing modality" problem, recent research has shifted toward flexible frameworks capable of handling arbitrary subsets of inputs. The pioneering work of HeMIS [havaei2016hemis] proposed learning a shared latent representation where first- and second-order statistics of available modalities are fused. This statistical fusion concept was extended by Dorent et al. [dorent2019hetero] in the U-HVED framework, which integrates a multi-modal VAE to reconstruct missing inputs. Other approaches employ knowledge distillation (KD) to transfer information from a "complete-modality" teacher to an "incomplete-modality" student [hu2020knowledge, wang2021acn]. However, KD methods typically require a complete set of modalities during the training phase, which may not always be available in retrospective cohorts. More recently, network architecture designs have moved toward adaptive feature fusion. RFNet [ding2021rfnet] utilizes region-aware fusion to selectively aggregate features based on image content. With the rise of Transformers, methods like mmFormer [zhang2022mmformer] and M2FTrans [m2ftrans] employ masked attention mechanisms to learn intra- and inter-modal correlations, effectively "masking out" missing tokens. M3AE [liu2023m3ae] adapts the Masked Autoencoder (MAE) paradigm to multi-modal settings. While these methods achieve state-of-the-art results, they generally treat modalities as independent, uncorrelated channels.
@@ 3.1 Problem Formulation
To address the challenges of incomplete modalities, we propose a method based on a decomposition of the Hounsfield Unit (HU) value. In multi-phase CT, the HU value at any spatial location $\mathbf{p}=(z,x,y)$ is a function of the time $\tau$ elapsed since the injection of a contrast agent. This function, $H(\mathbf{p},\tau)$, is known as the Time-Attenuation Curve (TAC) and describes the complete perfusion profile of the tissue at that location, as shown in Fig. 2. Clinically, a modality (e.g., Arterial, Venous, Delayed phase) corresponds to a single sample of this curve at a specific time $\tau$. A non-contrast CT represents the baseline at $\tau=0$.
@@ 3.2 Overall Architecture
Figure 3: (a) The overall architecture of TARDis. Input volumes are processed by a shared encoder. The features are projected and flattened into $K$ tokens. The modal-agnostic path (top) queries an embedding dictionary to extract consistent anatomic features. The modal-specific path (bottom) regresses a relative time $\tau$ and uses a Conditional Variational Autoencoder (CVAE) conditioned on the average anatomy and $\tau$ to sample dynamic features. Finally, the aggregated representations are assembled and decoded by a U-Net decoder. (b) The detailed workflow of the modal-specific path. Every token $X_{d,i}(\mathbf{p})$ contains the dynamic features of a neighboring subregion in different modalities. First, a $\tau_{i}$ is regressed based on dynamic features. Then we treat the time $\tau_{i}$ and modal-agnostic anatomic features $\hat{X}_{s}(\mathbf{p})$ as the conditions, and project the dynamic features of each position $\mathbf{p}$ into Gaussian means $\mu(\mathbf{p})$ and variances $\sigma^{2}(\mathbf{p})$. Then we reparameterize the tokens with $\epsilon(\mathbf{p})\in\mathcal{N}(0,I)$ and reconstruct the $\hat{X}_{d,i}(\mathbf{p})$. When inferring, we can sample $z(\mathbf{p})$ directly from $\mathcal{N}(0,I)$ and generate the corresponding dynamic representations with respect to the given $(\hat{X}_{s}(\mathbf{p}),\tau_{i})$
The overall architecture of TARDis is illustrated in Fig. 3(a). Unlike previous incomplete modality methods that use separate encoders for fixed inputs, we employ a shared encoder to extract features from a flexible number of modalities. When taking a series of registered incomplete multi-modal CT volumes $I_{N}$ as inputs, each of the input modality volumes $I_{i}\in\mathbb{R}^{1\times D\times H\times W},i\in\{1,2,\dots,N\}$ is independently processed by a shared 3D encoder, producing a feature map $F_{i}\in\mathbb{R}^{C\times D^{\prime}\times H^{\prime}\times W^{\prime}}$. These multi-modal features are then disentangled by a pair of projectors and flattened into sequences of tokens $X_{s}\in\mathbb{R}^{N\times C\times K}$ and $X_{d}\in\mathbb{R}^{N\times C\times K}$. Then the tokens are processed by two distinct paths. The modal-agnostic path extracts a common static anatomical representation $\hat{X}_{s}\in\mathbb{R}^{C\times K}$ using a shared embedding dictionary (Sec. 3.3). The modal-specific path estimates the phase time $\tau_{i}$ and uses a CVAE to model and sample the dynamic representations $\hat{X}_{d}\in\mathbb{R}^{N\times C\times K}$ for all modalities (Sec. 3.4). The derived anatomic representation and the multi-modal dynamic representations are adaptively assembled (Sec. 3.5), reshaped back to spatial dimensions, and passed to a decoder to predict the final segmentation mask. The entire network is trained jointly using the objectives detailed in Sec. 3.6.
@@ 3.3 Modal-Agnostic Representations
The static component represents the inherent anatomy, which should be consistent across all phases. However, features extracted from different contrast phases often contain phase-specific noise that obscures this underlying anatomy. Inspired by VQ-VAE [van2017neural], we construct a discrete learnable Embedding Dictionary $\mathcal{D}\in\mathbb{R}^{M\times C}$, containing $M$ anatomical vision queries, into which the anatomical information is distilled by quantization.
@@ 3.4 Sampling Based Modal-Specific Representations
The dynamic component varies significantly with time, reflecting the perfusion of contrast agents. To model this complex, time-dependent distribution, we design a generative module based on a CVAE. The details are shown in Fig. 3(b). This allows us to sample specific representations conditioned on the underlying anatomy and the particular time point.
@@
We model the distribution of continuous latent representations using a CVAE. The core idea is that dynamic enhancement is a function of both tissue type and time. Therefore, the generation of a specific dynamic feature $\hat{X}_{d}(\mathbf{p})$ at any position $\mathbf{p}$ is conditioned on two variables: the neighboring anatomic representation $\hat{X}_{s}(\mathbf{p})$, which provides the structural context, and the relative time $\tau_{i}$, which provides the phase context.
The CVAE’s objective is to learn a probabilistic mapping for the dynamic features, which consists of an encoder and a decoder. The encoder $q_{\psi}$ approximates the posterior Gaussian distribution $p(z(\mathbf{p})|X_{d}(\mathbf{p}),\hat{X}_{s}(\mathbf{p}),\tau)$:
@@
During training, this CVAE module is trained with KL Divergence and reconstruction loss. The total loss of the modal-specific path is:
@@
Crucially, the CVAE is conditioned on the output of the modal-agnostic module and the known phase. This aligns with our physical model where the dynamic enhancement $H_{d}$ depends on the static tissue $H_{s}$ and time $\tau$ in Eq. (2). By training this CVAE, it is capable of capturing the dynamic enhancement properties at position $\mathbf{p}$ for a specific phase $\tau$. This generative design provides a powerful mechanism for handling missing modalities. For flexible input modalities in the inference stage, a relative time $\tau_{i}$ can be predicted for each modality. And then a corresponding dynamic representation $\hat{X}_{d,i}$ can be generated by sampling equivalent $z_{i}$ from the prior distribution and decoding it with the inferred anatomy $\hat{X}_{s}$ and the target time $\tau_{i}$.

### v2

@@ Abstract
The accurate diagnosis and segmentation of tumors in contrast-enhanced Computed Tomography (CT) are fundamentally driven by the distinctive hemodynamic profiles of contrast agents over time. However, in real-world clinical practice, complete temporal dynamics are often hard to capture by strict radiation dose limits and inconsistent acquisition protocols across institutions, leading to a prevalent missing modality problem. Existing deep learning approaches typically treat missing phases as absent independent channels, ignoring the inherent temporal continuity of hemodynamics. In this work, we propose Time Attenuated Representation Disentanglement (TARDis), a novel physics-aware framework that redefines missing modalities as missing sample points on a continuous Time-Attenuation Curve. We first hypothesize that the latent feature can be disentangled into a time-invariant static component (anatomy) and a time-dependent dynamic component (perfusion). We achieve this via a dual-path architecture: a quantization-based path using a learnable embedding dictionary to extract consistent anatomical structures, and a probabilistic path using a Hemodynamic Conditional Variational Autoencoder to model dynamic enhancement conditioned on the estimated scan time. This design allows the network to infer missing hemodynamic features by sampling from the learned latent distribution. Extensive experiments on a large-scale multi-modal private abdominal CT dataset (2,282 patients) and two public datasets demonstrate that TARDis significantly outperforms state-of-the-art incomplete modality frameworks. Notably, our method maintains robust diagnostic performance even in extreme data-sparsity scenarios, highlighting its potential for reducing radiation exposure while maintaining diagnostic precision.
@@ 1 Introduction
Tumor segmentation in contrast-enhanced Computed Tomography (CT) is fundamentally dependent on the transient physiological dynamics of the contrast agent [bi2019artificial]. Crucially, the diagnostic boundaries of tumors are often defined by this time-dependent contrast-enhancement signal, which is unavailable in non-contrast acquisitions or incorrectly timed scans [bae2010intravenous]. Accurately capturing these dynamics usually requires a multi-phase acquisition protocol.
However, obtaining a standardized, complete set of phases is rarely feasible in real-world clinical practice. First, strict limits on radiation dose and scanning time often force clinicians to skip specific phases, creating a significant missing modality problem. Second, and perhaps more challenging, is the variability of acquisition protocols across different medical centers and tumor types. The definition of a specific phase often relies on fixed time delays that do not account for patient-specific cardiac output or variations in scanner hardware. As a result, a scan labeled as arterial at one hospital may effectively be a late arterial or early venous phase at another. This temporal mismatch introduces severe domain shifts, where the intensity distribution of the same modality varies drastically between datasets. Standard deep learning frameworks, which assume fixed, synchronized input channels, degrade severely when facing such inconsistent or incomplete data [javed2024robustness].
@@
Existing approaches to multi-modal segmentation typically rely on channel-concatenation strategies, which assume a fixed set of inputs. When a modality is missing, these models often fail or require retraining. To address this, recent incomplete modality frameworks have been proposed, utilizing techniques such as arithmetic mean fusion, knowledge distillation, or shared latent space mapping [dou2020unpaired, chen2019robust]. While these methods allow for flexible inputs, they generally treat different contrast phases as independent, uncorrelated channels or simple statistical variations. By ignoring the temporal dependency between phases, existing methods struggle to infer the missing dynamic information effectively, often resulting in suboptimal feature representation [shen2019brain, sharma2019missing].
@@
In this work, we propose Time Attenuated Representation Disentanglement (TARDis), a novel framework that addresses the incomplete modality challenge by modeling the underlying hemodynamics of CT. Our central hypothesis is that the Hounsfield Unit (HU) intensity can be explicitly decoupled into a time-invariant static component and a time-dependent dynamic component. Rather than treating missing modalities as empty data channels, we treat them as missing samples on the Time-Attenuation Curve (TAC). Our method employs a shared encoder to process a flexible number of input modalities. To learn robust representations, we introduce a dual-path disentanglement mechanism, as shown in Fig. 1. First, we extract a modal-agnostic anatomical representation using a learnable embedding dictionary, ensuring consistency across different scans. Second, we model the modal-specific dynamic representation by designing a physics-aware Conditional Variational Autoencoder (CVAE) [sohn2015learning], which is conditioned on the anatomy and the estimated scan time. This allows the network to probabilistically reconstruct the dynamic enhancement features of the tumor even when specific phases are missing. By explicitly disentangling anatomy from time-dependent perfusion, TARDis learns a physically grounded latent space that is robust to missing data, superior to standard fusion or masking strategies.
@@
Figure 3: (a) The overall architecture of TARDis. Input volumes are processed by a shared encoder. The features are projected and flattened into $K$ tokens. The modal-agnostic path (top) queries an embedding dictionary to extract consistent anatomic features. The modal-specific path (bottom) regresses a relative time $\tau$ and uses a Hemodynamic Conditional Variational Autoencoder (HCVAE) conditioned on the average anatomy and $\tau$ to sample dynamic features. Finally, the aggregated representations are assembled and decoded by a U-Net decoder. (b) The detailed workflow of the modal-specific path. Every token $X_{d,i}(\mathbf{p})$ contains the dynamic features of a neighboring subregion in different modalities. First, a $\tau_{i}$ is regressed based on dynamic features. Then we treat the time $\tau_{i}$ and modal-agnostic anatomic features $\hat{X}_{s}(\mathbf{p})$ as the conditions, and project the dynamic features of each position $\mathbf{p}$ into Gaussian means $\mu(\mathbf{p})$ and variances $\sigma^{2}(\mathbf{p})$. Then we reparameterize the tokens with $\epsilon(\mathbf{p})\in\mathcal{N}(0,I)$ and reconstruct the $\hat{X}_{d,i}(\mathbf{p})$. When inferring, we can sample $z(\mathbf{p})$ directly from $\mathcal{N}(0,I)$ and generate the corresponding dynamic representations with respect to the given $(\hat{X}_{s}(\mathbf{p}),\tau_{i})$
@@ 2.1 Multi-Modal Medical Image Segmentation
Multi-modal learning has become a cornerstone in medical image analysis, as combining complementary information from different modalities significantly improves diagnostic precision [9363915]. Early approaches focused on input-level fusion, where registered modalities are concatenated as multi-channel inputs to a Convolutional Neural Network [kamnitsas2017efficient, milletari2016v]. While effective, this assumes a fixed set of inputs, making the model fragile to missing data. Feature-level fusion strategies have subsequently emerged to better exploit cross-modal correlations. For instance, Dolz et al. [dolz2018hyperdense] proposed HyperDenseNet to facilitate dense connectivity across multi-modal streams. More recently, Transformer-based architectures have been adopted to model long-range dependencies across modalities. Hatamizadeh et al. [hatamizadeh2022unetr] introduced UNETR, which utilizes a vision transformer encoder to fuse volumetric data. Despite their success, these complete-modality frameworks rely on the assumption that all training and testing modalities are available. In clinical practice—particularly in multi-phase CT—scans are often incomplete due to radiation dose concerns or varying acquisition protocols, rendering these rigid fusion models ineffective without retraining or zero-filling strategies that often introduce artifacts.
@@ 2.2 Learning with Incomplete Modalities
To address the missing modality problem, recent research has shifted toward flexible frameworks capable of handling arbitrary subsets of inputs. The pioneering work of HeMIS [havaei2016hemis] proposed learning a shared latent representation where first- and second-order statistics of available modalities are fused. This statistical fusion concept was extended by Dorent et al. [dorent2019hetero] in the U-HVED framework, which integrates a multi-modal VAE to reconstruct missing inputs. Other approaches employ knowledge distillation (KD) to transfer information from a complete-modality teacher to an incomplete-modality student [hu2020knowledge, wang2021acn]. However, KD methods typically require a complete set of modalities during the training phase, which may not always be available in retrospective cohorts. More recently, network architecture designs have moved toward adaptive feature fusion. RFNet [ding2021rfnet] utilizes region-aware fusion to selectively aggregate features based on image content. With the rise of Transformers, methods like mmFormer [zhang2022mmformer] and M2FTrans [m2ftrans] employ masked attention mechanisms to learn intra- and inter-modal correlations, effectively masking out missing tokens. M3AE [liu2023m3ae] adapts the Masked Autoencoder (MAE) paradigm to multi-modal settings. While these methods achieve state-of-the-art results, they generally treat modalities as independent, uncorrelated channels.
@@ 3.1 Problem Formulation
To address the challenges of incomplete modalities, we propose a method based on a decomposition of the Hounsfield Unit (HU) value. In multi-phase CT, the HU value at any spatial location $\mathbf{p}=(z,x,y)$ is a function of the time $\tau$ elapsed since the injection of a contrast agent. This function, $H(\mathbf{p},\tau)$, is known as the Time-Attenuation Curve and describes the complete perfusion profile of the tissue at that location, as shown in Fig. 2. Clinically, a modality (e.g., Arterial, Venous, Delayed phase) corresponds to a single sample of this curve at a specific time $\tau$. A non-contrast CT represents the baseline at $\tau=0$.
@@ 3.2 Overall Architecture
The overall architecture of TARDis is illustrated in Fig. 3(a). Unlike previous incomplete modality methods that use separate encoders for fixed inputs, we employ a shared encoder to extract features from a flexible number of modalities. When taking a series of registered incomplete multi-modal CT volumes $I_{N}$ as inputs, each of the input modality volumes $I_{i}\in\mathbb{R}^{1\times D\times H\times W},i\in\{1,2,\dots,N\}$ is independently processed by a shared 3D encoder, producing a feature map $F_{i}\in\mathbb{R}^{C\times D^{\prime}\times H^{\prime}\times W^{\prime}}$. These multi-modal features are then disentangled by a pair of projectors and flattened into sequences of tokens $X_{s}\in\mathbb{R}^{N\times C\times K}$ and $X_{d}\in\mathbb{R}^{N\times C\times K}$. Then the tokens are processed by two distinct paths. The modal-agnostic path extracts a common static anatomical representation $\hat{X}_{s}\in\mathbb{R}^{C\times K}$ using a shared embedding dictionary (Sec. 3.3). The modal-specific path estimates the phase time $\tau_{i}$ and uses a Hemodynamic CVAE (HCVAE) to model and sample the dynamic representations $\hat{X}_{d}\in\mathbb{R}^{N\times C\times K}$ for all modalities (Sec. 3.4). The derived anatomic representation and the multi-modal dynamic representations are adaptively assembled (Sec. 3.5), reshaped back to spatial dimensions, and passed to a decoder to predict the final segmentation mask. The entire network is trained jointly using the objectives detailed in Sec. 3.6.
@@ 3.3 Modal-Agnostic Representations
The static component represents the inherent anatomy, which should be consistent across all phases. However, features extracted from different contrast phases often contain phase-specific noise that obscures this underlying anatomy. We employ VQ-VAE [van2017neural] for the anatomical branch because anatomy is deterministic and time-invariant. By forcing features to map to a finite, discrete Embedding Dictionary, the VQ mechanism acts as a robust information bottleneck. It effectively filters out phase-specific intensity variations and noise, ensuring that the extracted anatomical representation remains consistent regardless of whether the input is an arterial, venous, or non-contrast scan. We construct the learnable Embedding Dictionary $\mathcal{D}\in\mathbb{R}^{M\times C}$, containing $M$ anatomical vision queries, into which the anatomical information is distilled by quantization.
@@ 3.4 Sampling Based Modal-Specific Representations
The dynamic component varies significantly with time, reflecting the perfusion of contrast agents. To model this complex, time-dependent distribution, we design a generative module based on an HCVAE. The details are shown in Fig. 3(b). This allows us to sample specific representations conditioned on the underlying anatomy and the particular time point.
@@
We model the distribution of continuous latent representations using the designed HCVAE. The core idea is that dynamic enhancement is a function of both tissue type and time. Therefore, the generation of a specific dynamic feature $\hat{X}_{d}(\mathbf{p})$ at any position $\mathbf{p}$ is conditioned on two variables: the neighboring anatomic representation $\hat{X}_{s}(\mathbf{p})$, which provides the structural context, and the relative time $\tau_{i}$, which provides the phase context.
The HCVAE’s objective is to learn a probabilistic mapping for the dynamic features, which consists of an encoder and a decoder. The encoder $q_{\psi}$ approximates the posterior Gaussian distribution $p(z(\mathbf{p})|X_{d}(\mathbf{p}),\hat{X}_{s}(\mathbf{p}),\tau)$:
@@
During training, this HCVAE module is trained with KL Divergence and reconstruction loss. The total loss of the modal-specific path is:
@@
Crucially, the HCVAE is conditioned on the output of the modal-agnostic module and the known phase. This aligns with our physical model where the dynamic enhancement $H_{d}$ depends on the static tissue $H_{s}$ and time $\tau$ in Eq. (2). By training this HCVAE, it is capable of capturing the dynamic enhancement properties at position $\mathbf{p}$ for a specific phase $\tau$. This generative design provides a powerful mechanism for handling missing modalities. For flexible input modalities in the inference stage, a relative time $\tau_{i}$ can be predicted for each modality. And then a corresponding dynamic representation $\hat{X}_{d,i}$ can be generated by sampling equivalent $z_{i}$ from the prior distribution and decoding it with the inferred anatomy $\hat{X}_{s}$ and the target time $\tau_{i}$.